OpenStudy (anonymous):

Kellie called six friends with each call lasting 8 1/2 minutes. How long was Kellie on the phone?

OpenStudy (karolynxx):

51 minutes

OpenStudy (karolynxx):

= 6(8 1/2) = 6(17/2) = 102/2 or 51

OpenStudy (anonymous):

Okay. \[= 6(8 \frac{ 1 }{ 2 })\] = \[6(17/2) \] = 102/2 or 51 Answer (q#1): 51 minutes

OpenStudy (karolynxx):

kellie called six friends with each call lasting 8 and 1/2 ( mixed number) minutes how long was she on the phone. Ans: one call lasted for 8 and 1/2 mins = 8 mins 30 secs = 510 secs therefore, six calls lasted for = 6 X 510 secs = 3060secs therefore, in the whole day she was on the phone for 3060 secs = 51 minutes

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @tyaunasharae

OpenStudy (anonymous):

@karolynxx @Tom_Boy_Rebel Thankyou!